# On the SOS Rank of Simple and Diagonal Biquadratic Forms

### v2

@@ Abstract
We study the sum-of-squares (SOS) rank of simple and diagonal biquadratic forms. For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form. We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$, providing a general lower bound that extends the $3\times 3$ case. For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms. These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
@@ 1 Introduction
  • For any $m\geq 3$, there exists an $m\times m$ simple biquadratic form with SOS rank exactly $2m$, generalizing the $3\times 3$ case and providing a linear lower bound in $m$ (Theorem 2.7).
@@ 2.1 Extension to $m\times m$ Simple Forms
Theorem 2.7 immediately implies a lower bound for the maximal SOS rank over all $m\times m$ biquadratic forms:
$\mathrm{BSR}(m,m)\geq 2m\qquad(m\geq 3).$
Recall that the universal upper bound from [5] gives $\mathrm{BSR}(m,m)\leq m^{2}-1$. Thus the gap between the currently best lower and upper bounds grows quadratically with $m$. For $m=2$, it is known that $\mathrm{BSR}(2,2)=3$ (see [4]), which also equals 2m12m-1. It remains open whether $\mathrm{BSR}(m,m)$ can reach the order $O(m^{2})$ or remains linear in $m$.
@@ 5 Concluding Remarks
We have determined the exact maximum SOS rank ($6$) for $3\times 3$ simple biquadratic forms and proved an improved upper bound ($7$) for $3\times 3$ diagonal forms. Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$. These results demonstrate how structural constraints can either lower or raise the required number of squares in an SOS decomposition.

### v3

@@ Abstract
We study the sum-of-squares (SOS) rank of simple and diagonal biquadratic forms. For simple biquadratic forms in $3\times 3$ variables, we show that the maximum SOS rank is exactly $6$, attained by a specific six-term form. We further prove that for any $m\geq 3$, there exists an $m\times m$ simple biquadratic form whose SOS rank is exactly $2m$. Moreover, we show that for all $m,n\geq 3$, the maximum SOS rank over $m\times n$ simple biquadratic forms is at least $m+n$, which implies $\mathrm{BSR}(m,n)\geq m+n$. For diagonal biquadratic forms with nonnegative coefficients, we prove an SOS rank upper bound of $7$, improving the general bound of $8$ for $3\times 3$ forms. These results provide new lower and upper bounds on the worst-case SOS rank of biquadratic forms and highlight the role of structure in reducing the required number of squares.
@@ 1 Introduction
  • For any $m\geq 3$, there exists an $m\times m$ simple biquadratic form with SOS rank exactly $2m$, generalizing the $3\times 3$ case (Theorem 2.7).
  • For all $m,n\geq 3$, the maximum SOS rank of $m\times n$ simple biquadratic forms is at least $m+n$, implying $\mathrm{BSR}(m,n)\geq m+n$ (Theorem 2.10).
@@ 2.1 Extension to $m\times m$ Simple Forms
2.2 A General Lower Bound for $\mathrm{BSR}_{\mathrm{simple}}(m,n)$
We denote by $\mathrm{BSR}_{\mathrm{simple}}(m,n)$ the maximum SOS rank over all simple $m\times n$ biquadratic forms.
Lemma 2.8 (Monotonicity).
For any $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m,n+1)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
Proof.
Let $P({\bf x},{\bf y})$ be an $m\times n$ simple biquadratic form with $\operatorname{sos}(P)=\mathrm{BSR}_{\mathrm{simple}}(m,n)$. Consider the $m\times(n+1)$ simple biquadratic form
$P_{+}({\bf x},{\bf y},y_{n+1})=P({\bf x},{\bf y})+x_{1}^{2}y_{n+1}^{2}.$
Clearly, $\operatorname{sos}(P_{+})\leq\operatorname{sos}(P)+1=\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
We now show that $\operatorname{sos}(P_{+})\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$. Suppose, for contradiction, that $P_{+}$ admits an SOS decomposition with $R=\mathrm{BSR}_{\mathrm{simple}}(m,n)$ squares:
$P_{+}=\sum_{t=1}^{R}L_{t}({\bf x},{\bf y},y_{n+1})^{2},$
where each $L_{t}$ is bilinear in $\bf x$} and $({\bf y},y_{n+1})$. Write $L_{t}=M_{t}({\bf x},{\bf y})+N_{t}({\bf x})y_{n+1}$, where $M_{t}$ is bilinear in ${\bf x},{\bf y}$ and $N_{t}$ is linear in $\bf x$}. Expanding and comparing coefficients, we have
$\sum_{t=1}^{R}M_{t}({\bf x},{\bf y})^{2}=P({\bf x},{\bf y}),\qquad\sum_{t=1}^{R}N_{t}({\bf x})^{2}=x_{1}^{2},\qquad\sum_{t=1}^{R}M_{t}({\bf x},{\bf y})N_{t}({\bf x})=0.$
The first equality shows that the $M_{t}$ give an SOS representation of $P$ with $R$ squares. Since $R=\operatorname{sos}(P)$ is minimal, the $M_{t}$ are linearly independent as bilinear forms. The second equality implies that the $N_{t}$ are multiples of $x_{1}$: $N_{t}=c_{t}x_{1}$ with $\sum_{t}c_{t}^{2}=1$. Then the third equality becomes $\sum_{t}c_{t}M_{t}({\bf x},{\bf y})=0$, which is a linear dependence among the $M_{t}$, a contradiction. Hence, $\operatorname{sos}(P_{+})\geq R+1$, and so $\operatorname{sos}(P_{+})=\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$. Therefore, $\mathrm{BSR}_{\mathrm{simple}}(m,n+1)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$. ∎
By symmetry, the same argument applied to the variables $\bf x$} and $\bf y$} yields the following corollary.
Corollary 2.9 (Monotonicity in the first dimension).
For any $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m+1,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)+1$.
Theorem 2.10.
For all $m,n\geq 3$, we have $\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq m+n$. Consequently, $\mathrm{BSR}(m,n)\geq m+n$.
Proof.
We prove the inequality by induction on $k=m+n$.
Base case: $k=6$, i.e., $m=n=3$. By Theorem 2.5, $\mathrm{BSR}_{\mathrm{simple}}(3,3)=6=3+3$.
Inductive step: Assume the inequality holds for all $m^{\prime},n^{\prime}\geq 3$ with $m^{\prime}+n^{\prime}<k$. Consider $m,n\geq 3$ with $m+n=k$.
If $m>3$, then by Corollary 2.9,
$\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m-1,n)+1.$
Since (m1)+n=k1(m-1)+n=k-1, the induction hypothesis gives $\mathrm{BSR}_{\mathrm{simple}}(m-1,n)\geq(m-1)+n=m+n-1$. Hence,
$\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq(m+n-1)+1=m+n.$
If $m=3$, then $n=k-3\geq 3$. By Lemma 2.8,
$\mathrm{BSR}_{\mathrm{simple}}(3,n)\geq\mathrm{BSR}_{\mathrm{simple}}(3,n-1)+1.$
Since 3+(n1)=k13+(n-1)=k-1, the induction hypothesis gives $\mathrm{BSR}_{\mathrm{simple}}(3,n-1)\geq 3+(n-1)=n+2$. Hence,
$\mathrm{BSR}_{\mathrm{simple}}(3,n)\geq(n+2)+1=n+3=3+n.$
Thus, in both cases we have $\mathrm{BSR}_{\mathrm{simple}}(m,n)\geq m+n$.
Finally, since $\mathrm{BSR}(m,n)\geq\mathrm{BSR}_{\mathrm{simple}}(m,n)$, we obtain $\mathrm{BSR}(m,n)\geq m+n$ for all $m,n\geq 3$. ∎
Example 2.11 (Explicit family attaining the lower bound).
For any $m,n\geq 3$, define the simple biquadratic form
$Q_{m,n}({\bf x},{\bf y})=\sum_{i=1}^{m}x_{i}^{2}y_{i}^{2}+\sum_{i=1}^{m}x_{i}^{2}y_{i+1}^{2}+\sum_{j=m+1}^{n}x_{1}^{2}y_{j}^{2},$
where in the second sum we interpret $y_{m+1}=y_{1}$ when $m=n$, and otherwise $y_{m+1}$ is the $(m+1)$-th $y$ variable. Then $Q_{m,n}$ has exactly $m+n$ terms and satisfies $\operatorname{sos}(Q_{m,n})=m+n$.
Proof sketch.
The upper bound $\operatorname{sos}(Q_{m,n})\leq m+n$ is immediate since each term is already a square. For the lower bound, we show that any SOS decomposition of $Q_{m,n}$ must involve at least $m+n$ squares.
Consider the support set $S$ of $Q_{m,n}$, which consists of:
  • The diagonal pairs $(i,i)$ for $i=1,\dots,m$
  • The off-diagonal pairs $(i,i+1)$ for $i=1,\dots,m$ (with $y_{m+1}=y_{1}$ when $m=n$)
  • The pairs $(1,j)$ for $j=m+1,\dots,n$ (when $n>m$)
This gives exactly $|S|=m+n$ terms.
Assume $Q_{m,n}=\sum_{t=1}^{R}L_{t}({\bf x},{\bf y})^{2}$ with each $L_{t}$ bilinear. By comparing coefficients of $x_{i}^{2}y_{j}^{2}$, we find that for each $(i,j)\in S$, the vectors $\mathbf{C}_{ij}=(c_{ij}^{(1)},\dots,c_{ij}^{(R)})$ satisfy $\|\mathbf{C}_{ij}\|^{2}=1$, where $c_{ij}^{(t)}$ are the coefficients of $L_{t}$. Moreover, for any two distinct pairs $(i,j),(p,q)\in S$, the cross-term conditions force $\mathbf{C}_{ij}\cdot\mathbf{C}_{pq}=0$. Thus the vectors $\{\mathbf{C}_{ij}:(i,j)\in S\}$ form an orthonormal set in $\mathbb{R}^{R}$, implying $R\geq|S|=m+n$. ∎
@@ 2.2 A General Lower Bound for $\mathrm{BSR}_{\mathrm{simple}}(m,n)$
Theorem 2.7 exhibits an explicit $m\times m$ simple biquadratic form with SOS rank exactly $2m$, showing that the lower bound in Theorem 2.10 is tight for $m=n$. For non-square cases, the question of whether $\mathrm{BSR}_{\mathrm{simple}}(m,n)$ can exceed $m+n$ remains open. Theorems 2.7 and 2.10 together imply
$\mathrm{BSR}(m,n)\geq m+n\qquad(m,n\geq 3).$
Recall that the universal upper bound from [5] gives $\mathrm{BSR}(m,n)\leq mn-1$. For the square case $m=n$, this leaves a gap between $2m$ and $m^{2}-1$, which grows quadratically with $m$. For $m=2$, it is known that $\mathrm{BSR}(2,2)=3$ (see [4]), which equals 2m12m-1. It remains open whether $\mathrm{BSR}(m,m)$ can reach the order $O(m^{2})$ or remains linear in $m$.
@@ 5 Concluding Remarks
We have determined the exact maximum SOS rank ($6$) for $3\times 3$ simple biquadratic forms and proved an improved upper bound ($7$) for $3\times 3$ diagonal forms. Moreover, we have shown that for every $m\geq 3$ there exists an $m\times m$ simple biquadratic form whose SOS rank equals $2m$, extending the $3\times 3$ case and providing a linear lower bound in $m$. Furthermore, we proved that for all $m,n\geq 3$, the maximum SOS rank of $m\times n$ simple biquadratic forms is at least $m+n$, which implies $\mathrm{BSR}(m,n)\geq m+n$.
Table 1 summarizes the current knowledge on the worst-case SOS rank $\mathrm{BSR}(m,n)$ for small values of $m$ and $n$.
Table 1: Known bounds on $\mathrm{BSR}(m,n)$
$(m,n)$ Lower bound Upper bound
$(2,2)$ 33 (exact) 33
$(3,2)$ 44 (exact) 44
$(m,2)$, $m\geq 4$ $m+1$ 2m12m-1
$(3,3)$ 66 88 (conjectured $6$)
$(m,n)$, $\geq 3 m+nm+n$ mn1mn-1
@@
For larger $m$ and $n$, the gap between the lower bound $m+n$ and the general upper bound mn1mn-1 widens. Determining the true growth rate of $\mathrm{BSR}(m,n)$ whether it remains linear or can be quadratic in $m$ and $n$ is a challenging direction for future research.